OpenStudy (anonymous):

enter the solution of the system of equations below as an ordered pair. 4x=3y+20 -20y=-20x+100

OpenStudy (ajprincess):

-20y=-20x+100 20x=20y+100 x=y+5-(1) 4x=3y+20-(2) (1)*4=>4x=4y+20-(3) (3)-(2)=>0=y substitute this in (1) to find x.

OpenStudy (anonymous):

your answer is (35,40) where x=35 and y=40 please if my answer is correct then give me a medal thanks!

OpenStudy (anonymous):

i dont think it works for the bottom equation

OpenStudy (anonymous):

bro both sides are equal by these values.

OpenStudy (ajprincess):

yes u r rigth. it dusn't work for the sec eqn.

OpenStudy (anonymous):

you get -800=-600.so whats the right answer?

OpenStudy (ajprincess):

chck mine after solvng for x.

OpenStudy (anonymous):

is yours saying the value of x=1?

OpenStudy (anonymous):

ok i will solve it again!

OpenStudy (anonymous):

okay thank you!

OpenStudy (ajprincess):